OpenStudy (anonymous):

If 15m = 63, find the value of 4m.

OpenStudy (anonymous):

m = 63/15 4m = 63/15 * 4

hero (hero):

4m = 84/5

hero (hero):

or 16.8

OpenStudy (anonymous):

right

OpenStudy (anonymous):

84/5